# Dynamical phase transitions in the non-reciprocal Ising model

### v1

@@ Abstract
Non-reciprocal interactions in many-body systems lead to time-dependent states, commonly observed in biological, chemical, and ecological systems. The stability of these states in the thermodynamic limit, as well as the criticality of the phase transition from static to time-dependent states remains an open question. To tackle these questions, we study a minimalistic system endowed with non-reciprocal interactions: an Ising model with two spin species having opposing goals. The mean-field equation predicts three stable phases: disordered, ordered, and a time-dependent swap phase. Large scale numerical simulations support the following: (i) in 2D, the swap phase is destabilized by defects; (ii) in 3D, the swap phase is stable, and has the properties of a time-crystal; (iii) the transition from disorder to swap in 3D is characterized by the critical exponents of the 3D XY model, in agreement with the emerging continuous symmetry of time translation invariance; (iv) when the two species have fully anti-symmetric couplings, the static-order phase is unstable in any dimension due to droplet growth; (v) in the general case of asymmetric couplings, static order can be restored by a droplet-capture mechanism preventing the droplets from growing indefinitely. We provide details on the full phase diagram which includes first- and second-order-like phase transitions and study the coarsening dynamics of the swap as well as the static-order phases.
@@ I Introduction
Many-body systems with non-reciprocal interactions emerge in contexts ranging from neuroscience [1, 2, 3] and social networks [4, 5, 6] to ecology [7, 8, 9, 10] and open quantum systems [11, 12, 13]. These systems exhibit rich collective behavior including time-dependent states that can arise because the dynamics does not follow from the minimisation of a single potential function [14, 15, 16, 17, 18, 19, 20, 21]. However, it remains unknown whether these time-dependent states can be thought of as proper non-equilibrium phases of matter, in the sense of the term used in statistical physics; and if so, whether they exhibit well-defined phase transitions and critical behavior. In particular, we are interested in the fate of these time-dependent states in systems made of locally coupled units subject to thermal noise in finite spatial dimensions when they are taken to the thermodynamic limit.
In order to address this question, we consider a non-reciprocal version of the Ising model, in which two species of spins have incompatible goals: spins of species A tend to align with spins of species B, while spins B tend to anti-align with spins A. This leads to a cyclic dynamics where spins of each species tend to flip in turn. We study this system through analytical arguments and extensive Monte-Carlo simulations. A short account of our results is presented in Letter form in Ref. [22]; we summarize here the most salient points. First, numerical simulations indicate that in our model, a time-dependent phase with long-range spatial and temporal order indeed exists in three dimensions. In this phase, the magnetization oscillates for arbitrarily long times. The system acts as a perfect clock in the thermodynamic limit, in the sense that the coherence time of the oscillations diverges with the size of the system. Second, a continuous phase transition between this oscillatory phase and a disordered (i.e. paramagnetic) phase, corresponding to a Hopf bifurcation in mean-field, is observed. Numerical evidence shows that it is associated with scale invariance, and the corresponding static critical exponents are compatible with the 3D XY universality class. Third, we found that a static ordered phase (similar to a ferromagnetic phase) can exist when non-reciprocal interactions are present, but only if there is some symmetric coupling between the species. However, we found no evidence of a direct transition between the oscillatory phase and the static-order phase, in contrast with mean-field predictions.
@@ II Background
II.1 Non-reciprocal interactions between groups
There is a rich literature on nonequilibrium lattice models, including variants of the Ising model with asymmetric couplings exhibiting a variety of complex phenomena [23, 24, 25, 26, 27]. A distinguishing feature of the non-reciprocal Ising model introduced in this paper is the presence of non-reciprocal interactions between two groups of spins of comparable sizes. This is a rather generic feature that arises in a variety of physical contexts, some of which can be captured by Ising or Ising-like minimal models. Examples include:
  • Nonlinear parametric resonators, that can be mapped to Ising spins. Networks of coupled parametric oscillators have been used as platforms to realize (reciprocal) “Ising machines” that solve combinatorial optimization problems [28, 29, 30, 31, 32, 29, 33, 34]. This is essentially because a single of these resonators implements a double-well potential, and a linear coupling effectively produces a (reciprocal) Ising-like coupling. Reference [35] shows experimentally and analytically that driven non-identical parametric resonators can be mapped into non-reciprocally coupled Ising spins. A network of such resonators could directly realize the NR Ising model discussed in the present work, as well as implement models of learning with non-reciprocal couplings between the artifical neurons [36, 3, 2, 37, 38, 39, 40].
  • Models of collective opinion dynamics, such as in the voter model [41, 23], where the values of the spins represent different opinions. Different communities may have asymmetric interactions: for instance, this is the case of conformists and contrarians (or alternatively trend-setters) [42, 43, 41, 44, 45, 46, 47, 48, 49]. Similarly, spin models are used to model game theory and decision making of economic agents [6].
  • Models of the collective response of transmembrane receptors, that have been put forward for the high sensitivity of cells to external chemicals in chemotaxis [50, 51, 52, 53, 54]. It has been argued that non-reciprocal couplings between the receptors and the kinases that serve as an output signal is required to account for experimental observations [55, 56, 57].
@@ II.1 Non-reciprocal interactions between groups
  • Models of spinor Bose-Einstein condensates based on two populations of Ising spins in a cavity (a Dicke model) that effectively interact non-reciprocally have been considered [11, 66, 67] to account for non-stationary dynamics observed in experiments [68].
  • Model inference in which symmetric and asymmetric Ising models have also been used as a variational ansatz [69, 70, 71, 72, 37, 73].
In several of these classes of systems, it has been conjectured that criticality may play a nontrivial role [74, 75, 76, 77, 78, 79, 80]. Our non-reciprocal Ising model can be seen as a stripped down version of these situations. In realistic cases, (i) there may be more than two populations (ii) the spins may take more than two possible states and (iii) the network topology may be more complex than $\mathbb{Z}^{d}$. While these differences are likely to induce additional complexities and phenomena, we expect that the basic features of the non-reciprocal Ising model will persist for systems in the same universality class [23, 75]. In particular, we expect that locally-coupled many-body systems undergoing a Hopf bifurcation in mean-field will enter this universality class even if they are not originally composed of coupled spin-like degrees of freedom.
@@ II.2 Phase transitions to time-dependent states
An underlying motivation for this work is to construct a bridge between bifurcations in non-variational systems and phase transitions out of equilibrium. In the equilibrium Ising model, for instance, the ferromagnet/paramagnet transition corresponds to a pitchfork bifurcation in mean-field. This means that the scalar order parameter $\phi$ (e.g. the magnetization) follows the dynamics $(d/dt)\phi=\alpha\phi-\phi^{3}$ in which $\alpha\propto T-T_{\text{c}}$ is proportional to the distance to the critical temperature. Conversely, the thermal field theory corresponding to locally-coupled pitchfork bifurcations where
$\partial_{t}\phi=\alpha\phi-\phi^{4}+D\Delta\phi+\eta(t,r),$ (1)
@@
In non-reciprocal and other non-equilibrium systems, the dynamics is typically non-variational,
$\Gamma\partial_{t}\phi=f[\phi]\neq-\frac{\delta\mathcal{F}}{\delta\phi}$ (3)
so bifurcations with no analogue at equilibrium can occur, in particular those leading to limit cycles. The simplest instance is the Hopf bifurcation, which describes the appearance of a limit cycle from a fixed point (unless otherwise specified, “fixed points” in this paper refer to fixed points of the physical dynamics, not fixed points of a renormalization group (RG) flow), see Refs. [82, 83]. This bifurcation is described by the equation
@@
for the complex number $z$ and takes place at $\alpha=0$. When $z$ is replaced by a complex field $\psi$ with appropriate diffusive couplings, the corresponding equation is known as the complex Ginzburg-Landau equation [84]. The fate of this bifurcation has been analyzed using perturbative RG in Refs. [85, 86], and later reproduced and generalized in Refs. [87, 88]. In these works, a phase transition generalizing the Hopf bifurcation is predicted based on a $d=4-\epsilon$ expansion and formally related to the RG fixed point of model A dynamics with $O(2)$ symmetry (i.e., to the XY universality class). Using this analogy Refs. [85, 86] predict that this phase transition, which leads to an oscillating phase with long-range order, should exist for space dimension $d>2$, while no coherent oscillations are expected in $d<2$, and quasi-long-range order in conjectured in $d=2$ 111 References [85, 86] conjecture that quasi-long-range may be present at $d=2$, but more recent studies based on a mapping to a version of the Kardar-Parisi-Zhang (KPZ) equation with compact variables suggest that the phase with algebraic order does not persist in the limit of large systems (nor an equivalent of the Berezinskii-Kosterlitz-Thouless (BKT) transition) but are replaced with a stretched exponential decay of correlations corresponding to the rough phase of the KPZ equation [101, 102]. In addition, Refs. [156] argue that the modified effective interaction between spiral defects (compared to that between XY vortices) allows for a finite density of defects at any temperature, potentially suppressing the algebraic order. . To the best of our knowledge, these predictions have only been tested qualitatively, in Refs. [90, 91], where numerical simulations of coupled three-state oscillators were performed for lattices of size $L^{d}$ going up to $L=80$ in $d=3$. A comparison of the rescaled finite-size scaling data of Refs. [90, 91] with Ising and XY exponents suggested a better collapse with XY exponents. The results of our large-scale simulations with lattices of sizes up to $L=320$ in $d=3$ are in agreement with the predictions of Refs. [85, 86] and with the qualitative observations of Refs. [90, 91]. In particular, we obtain values of the critical exponents that are compatible with XY exponents but not with Ising exponents, within one standard deviation.
@@
The Hopf bifurcation is not the only way to produce a limit cycle. Another bifurcation relevant here is the infinite-period bifurcation (or SNIC bifurcation) that describes the transition from a static ordered (symmetry-breaking) state to a limit cycle [82]. To the best of our knowledge, the field-theoretical version of this bifurcation has not been studied. It has been conjectured that it does not exist in spatially extended systems of arbitrary dimension in Ref. [92] based on the instability of the static ordered state in the three-state oscillator model of Refs. [90, 91]. Our simulations suggest that a static ordered state can exist despite the presence of non-reciprocal coupling (provided that they are not fully anti-symmetric, at least in our model), but even in this case we did not find evidence of any direct transition between the static order and oscillating phases.
@@ II.3 Stability of spatially coherent oscillations
Our work is also directly related to the question of the stability and the coherence of time-periodic phases in noisy locally-coupled classical many-body systems. By considering the nucleation and growth of droplets from fluctuations, Ref. [93, 94, 95, 96] argued that coherent oscillations are unstable in the thermodynamic limit if they arise from the spontaneous breaking of a discrete time-translation invariance (in particular when the period of oscillation is commensurate with the discrete time unit characterizing the evolution)  [^2] Note that these arguments require that the system is isotropic and that the noise is unbounded so that it can create arbitrarily large droplets [95, 96].. This is not the case in our model, where the oscillating phase arises from the spontaneous breaking of a continuous time-translation invariance (and the period is generically incommensurate). The case of spontaneously broken continuous time-translation invariance is considered in Refs. [95, 98, 96] using a mapping between the desynchronization of local oscillators and the roughening of surfaces described by the Kardar-Parisi-Zhang (KPZ) equation [99]. Based on this mapping, it is argued that temporal order can exist in $d>2$ but not in $d\leq 2$ (in isotropic systems). Note that in the mapping to the KPZ equation, the angular nature of phase variables (defined on a compact space) is neglected, which amounts to ignoring topological defects, expected not to affect the large-scale physics [98]. A compact version of KPZ equation has been studied in Refs. [100, 101, 102, 103], but not in isotropic systems in $d=3$. Similarly, Refs. [104, 87] suggested an analogy with the Mermin-Wagner theorem [105, 81] in the time domain, according to which the fluctuations of the Goldstone modes associated with broken time-translation invariance destroy order in $d\leq 2$.
References [106, 107, 108, 109, 110, 111, 112, 113] performed numerical simulations of cellular automata and coupled map lattices (see Ref. [107] for an introduction) in various space dimensions, from which it emerged that periodic oscillations are possible for $d\geq 3$ within this class of sytems [108, 114, 115]. Similar behaviors were reported in simulations of predator-prey models on lattice [116, 117, 118, 119, 120, 121, 23, 27]. These results are in agreement with the theoretical arguments reviewed above, and are are also consistent with our numerical observations. Similar questions arise in the study of the synchronization of locally coupled oscillators with randomly distributed frequencies (like in the Kuramoto model), see § IV.A of Ref. [122] and references therein, but the physics is very different because of the random frequencies.

### v2

# Dynamical phase transitions in the nonreciprocal Ising model

@@ Abstract
Nonreciprocal interactions in many-body systems lead to time-dependent states, commonly observed in biological, chemical, and ecological systems. The stability of these states in the thermodynamic limit and the critical behavior of the phase transition from static to time-dependent states are not yet fully understood. To address these questions, we study a minimalistic system endowed with nonreciprocal interactions: an Ising model with two spin species having opposing goals. The mean-field equation predicts three stable phases: disorder, static order, and a time-dependent swap phase. Large scale numerical simulations support the following: (i) in 2D, the swap phase is destabilized by defects; (ii) in 3D, the swap phase is stable, and has the properties of a time crystal; (iii) the transition from disorder to swap in 3D is characterized by the critical exponents of the 3D XY model, and corresponds to the breaking of a continuous symmetry, time translation invariance; (iv) when the two species have fully anti-symmetric couplings, the static-order phase is unstable in any finite dimension due to droplet growth; (v) in the general case of asymmetric couplings, static order can be restored by a droplet-capture mechanism preventing the droplets from growing indefinitely. We provide details on the full phase diagram which includes first- and second-order-like phase transitions and study how the system coarsens into swap and static-order states.
@@ I Introduction
Many-body systems with nonreciprocal interactions emerge in contexts ranging from neuroscience [1, 2, 3] and social networks [4, 5, 6] to ecology [7, 8, 9, 10] and open quantum systems [11, 12, 13]. These systems exhibit rich collective behavior including time-dependent states that can arise because the dynamics does not follow from the minimization of a single potential function [14, 15, 16, 17, 18, 19, 20, 21]. However, it remains unknown whether these time-dependent states can be thought of as proper non-equilibrium phases of matter, in the sense of the term used in statistical physics; and if so, whether they exhibit well-defined phase transitions and critical behavior. In particular, we are interested in the fate of these time-dependent states in systems made of locally coupled units subject to thermal noise in finite spatial dimensions when they are taken to the thermodynamic limit.
In order to address this question, we consider a nonreciprocal version of the Ising model, in which two species of spins have incompatible goals: spins of species A tend to align with spins of species B, while spins B tend to anti-align with spins A. This leads to a cyclic dynamics where spins of each species tend to flip in turn. We study this system through analytical arguments and extensive Monte-Carlo simulations. A short account of our results is presented in Letter form in Ref. [22]; we summarize here the most salient points. First, numerical simulations indicate that in our model, a time-dependent phase with long-range spatial and temporal order indeed exists in three spatial dimensions. In this phase, the magnetization oscillates for arbitrarily long times. The system acts as a perfect clock in the thermodynamic limit, in the sense that the coherence time of the oscillations diverges with the size of the system. Second, a continuous phase transition between this oscillatory phase and a disordered (i.e. paramagnetic) phase, corresponding to a Hopf bifurcation in mean-field, is observed. Numerical evidence shows that it is associated with scale invariance, and the corresponding static critical exponents are compatible with the 3D XY universality class. Third, we found that a static ordered phase (similar to a ferromagnetic phase) can exist when nonreciprocal interactions are present, but only if there is some symmetric coupling between the species. However, we found no evidence of a direct transition between the oscillatory phase and the static-order phase, in contrast with mean-field predictions.
@@ II Background
The following section provides background and context on different aspects related to our work, as well as relations to existing literature. It can be skipped on first reading.
II.1 Nonreciprocal interactions between groups
There is a rich literature on nonequilibrium lattice models, including variants of the Ising model with asymmetric couplings exhibiting a variety of complex phenomena [23, 24, 25, 26, 27]. A distinguishing feature of the nonreciprocal Ising model introduced in this paper is the presence of macroscopic nonreciprocal interactions between two groups of spins of comparable sizes. This is a rather generic feature that arises in a variety of physical contexts, some of which can be captured by Ising or Ising-like minimal models. Examples include:
  • Nonlinear parametric resonators, that can be mapped to Ising spins. Networks of coupled parametric oscillators have been used as platforms to realize (reciprocal) “Ising machines” that solve combinatorial optimization problems [28, 29, 30, 31, 32, 29, 33, 34]. This is essentially because a single of these resonators implements a double-well potential, and a linear coupling effectively produces a (reciprocal) Ising-like coupling. Reference [35] shows experimentally and analytically that driven non-identical parametric resonators can be mapped into nonreciprocally coupled Ising spins. A network of such resonators could directly realize the nonreciprocal Ising model discussed in the present work, as well as implement models of learning with nonreciprocal couplings between the artifical neurons [36, 3, 2, 37, 38, 39, 40].
  • Models of collective opinion dynamics, such as in the voter model [41, 23], where the values of the spins represent different opinions. Different communities may have asymmetric interactions: for instance, this is the case of conformists and contrarians (or alternatively trend-setters) [42, 43, 41, 44, 45, 46, 47, 48, 49]. Similarly, spin models have been used to model game theory and decision making of economic agents [6].
  • Models of the collective response of transmembrane receptors, that have been put forward for the high sensitivity of cells to external chemicals in chemotaxis [50, 51, 52, 53, 54]. It has been argued that nonreciprocal couplings between the receptors and the kinases that serve as an output signal is required to account for experimental observations [55, 56, 57].
@@ II.1 Nonreciprocal interactions between groups
  • Models of spinor Bose-Einstein condensates based on two populations of Ising spins in a cavity (a Dicke model) that effectively interact nonreciprocally have been considered [11, 66, 67] to account for non-stationary dynamics observed in experiments [68].
  • Model inference in which symmetric and asymmetric Ising models have also been used as a variational ansatz [69, 70, 71, 37, 72].
In several of these classes of systems, it has been conjectured that criticality may play a nontrivial role [73, 74, 75, 76, 77, 78, 79]. Our nonreciprocal Ising model can be seen as a stripped down version of these situations. In realistic cases, (i) there may be more than two populations (ii) the spins may take more than two possible states and (iii) the network topology may be more complex than $\mathbb{Z}^{d}$ (see e.g. Refs. [80, 81]). While these differences are likely to induce additional complexities and phenomena, we expect that the basic features of the nonreciprocal Ising model will persist for systems in the same universality class [23, 74]. In particular, we expect that locally-coupled many-body systems undergoing a Hopf bifurcation in mean-field will enter this universality class even if they are not originally composed of coupled spin-like degrees of freedom.
@@ II.2 Phase transitions to time-dependent states
An underlying motivation for this work is to construct a bridge between bifurcations in non-variational systems and phase transitions out of equilibrium. In the equilibrium Ising model, for instance, the ferromagnet/paramagnet transition corresponds to a pitchfork bifurcation in mean-field. This means that the scalar order parameter $\phi$ (e.g. the magnetization) follows the dynamics $({\rm d}/{\rm d}t)\phi=\alpha\phi-\phi^{3}$ in which $\alpha\propto T-T_{\text{c}}$ is proportional to the distance from the critical temperature. Conversely, the thermal field theory corresponding to locally-coupled pitchfork bifurcations where
$\partial_{t}\phi=\alpha\phi-\phi^{3}+D\Delta\phi+\eta(t,r),$ (1)
@@
In nonreciprocal and other non-equilibrium systems, the dynamics of the (not necessarily scalar) order parameter $\vec{\phi}$ is typically non-variational,
$\Gamma\partial_{t}\vec{\phi}=\vec{f}[\vec{\phi}]\neq-\frac{\delta\mathcal{F}}{\delta\vec{\phi}}$ (3)
so bifurcations with no analogue at equilibrium can occur, in particular those leading to limit cycles. The simplest instance is the Hopf bifurcation, which describes the appearance of a limit cycle from a fixed point, see Refs. [83, 84]. Unless otherwise specified, “fixed points” in this paper refer to fixed points of the physical dynamics, not fixed points of a renormalization group (RG) flow. The Hopf bifurcation describes a variety of spatially extended systems [85] ranging from oscillatory chemical reactions and extended lasers [86] to models of mammalian hearing [87, 88]. This bifurcation is described by the equation
@@
for the complex numbers $z$ and takes place at $\alpha=0$. When $z$ is replaced by a complex field $\psi$ with appropriate diffusive couplings, the corresponding equation is known as the complex Ginzburg-Landau equation [86]. The fate of this bifurcation has been analyzed using perturbative RG in Refs. [89, 90], and later reproduced and generalized in Refs. [91, 92]. In these works, a phase transition generalizing the Hopf bifurcation is predicted based on a $d=4-\epsilon$ expansion and formally related to the RG fixed point of model A dynamics with $O(2)$ symmetry (i.e., to the XY universality class). Using this analogy Refs. [89, 90] predict that this phase transition, which leads to an oscillating phase with long-range order, should exist for space dimension $d>2$, while no coherent oscillations are expected in $d<2$, and quasi-long-range order in conjectured in $d=2$ 111 References [89, 90] conjecture that quasi-long-range may be present at $d=2$, but more recent studies based on a mapping to a version of the Kardar-Parisi-Zhang (KPZ) equation with compact variables suggest that the phase with algebraic order does not persist in the limit of large systems (nor an equivalent of the Berezinskii-Kosterlitz-Thouless (BKT) transition) but are replaced with a stretched exponential decay of correlations corresponding to the rough phase of the KPZ equation [106, 107]. In addition, Refs. [177] argue that the modified effective interaction between spiral defects (compared to that between XY vortices) allows for a finite density of defects at any temperature, potentially suppressing the algebraic order.. To the best of our knowledge, these predictions have only been tested qualitatively, in Refs. [94, 95], where numerical simulations of coupled three-state oscillators were performed for lattices of size $L^{d}$ going up to $L=80$ in $d=3$. A comparison of the rescaled finite-size scaling data of Refs. [94, 95] with Ising and XY exponents suggested a better collapse with XY exponents. The results of our large-scale simulations with lattices of sizes up to $L=320$ in $d=3$ are in agreement with the predictions of Refs. [89, 90] and with the qualitative observations of Refs. [94, 95]. In particular, we obtain values of the critical exponents that are in very good agreement with XY exponents but not with Ising exponents.
@@
The Hopf bifurcation is not the only way to produce a limit cycle. Another relevant bifurcation is the infinite-period bifurcation (or SNIC bifurcation) that describes the creation of a limit cycle by a saddle-node bifurcation that connects heteroclinic orbits together [83]. Here, multiple SNIC bifurcations (related by symmetry) arise at the same time in the mean-field transition from a static ordered (symmetry-breaking) state to a limit cycle. To the best of our knowledge, the field-theoretical version of this class of bifurcations has not been studied. It has been conjectured that it does not exist in spatially extended systems of arbitrary dimension in Ref. [96] based on the instability of the static ordered state in the three-state oscillator model of Refs. [94, 95]. Our simulations suggest that a static ordered state can exist despite the presence of nonreciprocal coupling (provided that they are not fully anti-symmetric, at least in our model), but even in this case we did not find evidence of any direct transition between the static-order and oscillating phases.
@@ II.3 Stability of spatially coherent oscillations
Our work is also directly related to the question of the stability and spatial coherence of time-periodic phases in noisy locally-coupled classical many-body systems. By considering the nucleation and growth of droplets from fluctuations, Ref. [97, 98, 99, 100] argued that spatially coherent oscillations are unstable in the thermodynamic limit if they arise from the spontaneous breaking of a discrete time-translation invariance (in particular when the period of oscillation is commensurate with the discrete time unit characterizing the evolution)  [^2] Note that these arguments require that the system is isotropic and that the noise is unbounded so that it can create arbitrarily large droplets [99, 100].. This is not the case in our model, where the oscillating phase arises from the spontaneous breaking of a continuous time-translation invariance (and the period is generically incommensurate with the discrete time unit). The case of spontaneously broken continuous time-translation invariance is considered in Refs. [99, 102, 100] using a mapping between the desynchronization of local oscillators and the roughening of surfaces described by the Kardar-Parisi-Zhang (KPZ) equation [103]. In the KPZ universality class, surfaces are always rough in $d\leq 2$, while two phases (rough or smooth) are possible for $d>2$, see e.g. § 5.7.4 of Ref. [104]. Based on this mapping, it is argued that temporal order can exist in $d>2$ but not in $d\leq 2$ in isotropic systems. Note that in the mapping to the KPZ equation, the angular nature of phase variables (defined on a compact space) is neglected, which amounts to ignoring topological defects, expected not to affect the large-scale physics [102]. A compact version of the KPZ equation has been studied in Refs. [105, 106, 107, 108], but not in isotropic systems in $d=3$. Similarly, Refs. [109, 91] suggested an analogy with the Mermin-Wagner theorem [110, 82] in the time domain, according to which the fluctuations of the Goldstone modes associated with broken time-translation invariance destroy order in $d\leq 2$.
References [111, 112, 113, 114, 115, 116, 117, 118] performed numerical simulations of cellular automata and coupled map lattices (see Ref. [112] for an introduction) in various space dimensions, from which it emerged that periodic oscillations are possible for $d\geq 3$ within this class of sytems [113, 119, 120]. Similar behaviors were reported in simulations of predator-prey models on lattice [121, 122, 123, 124, 125, 126, 23, 27]. These results are in agreement with the theoretical arguments reviewed above, and are are also consistent with our numerical observations. Similar questions arise in the study of the synchronization of locally coupled oscillators with randomly distributed frequencies (like in the Kuramoto model), see § IV.A of Ref. [127] and references therein, but the physics is different because of the random frequencies.